OpenStudy (anonymous):

(6x+1)^2 +8(6x+1) + 16=0 Help me lol..

myininaya (myininaya):

u=6x+1 u^2=(6x+1)^2 \[u^2+8u+16=0\] \[(u+4)(u+4)=0\] \[(u+4)^2=0\] \[u+4=0\] \[u=-4\] but remember u=6x+1 so 6x+1=-4 6x=-4-1 6x=-5 x=-5/6

OpenStudy (anonymous):

YOU ARE AWESOME♥♥♥♥

myininaya (myininaya):

:)

myininaya (myininaya):

if you like my answer give me a medal :)